OpenStudy (anonymous):

Given below is the odd-number function. Which of the following are equal to O(7)? THERE ARE MULTIPLE RIGHT ANSWERS LIST THEM ALL PLEASE!!! O(n) = 2n - 1 A. The seventh odd number B. 13 C. 7 D. O(6) + 2 Given below is the odd-number function. Which of the following are equal to O(7)? THERE ARE MULTIPLE RIGHT ANSWERS LIST THEM ALL PLEASE!!! O(n) = 2n - 1 A. The seventh odd number B. 13 C. 7 D. O(6) + 2 @Mathematics

OpenStudy (anonymous):

Are you asking these questions for help or to see if people answer correctly?

OpenStudy (anonymous):

i need help.

OpenStudy (anonymous):

I take online courses and sometimes I can't figure out what to do so I post on here in hopes of getting help.

OpenStudy (anonymous):

Cool cool, just checking - it's annoying to spend time giving people answers who are only here to test others ;) Genuine help is what this place is for though. O(n) = 2n - 1 O(7) = 2*7 - 1 = 13 A. Seventh odd number = 13. Yes. B. 13 = 13. Yes. C. 7. No. D. O(6)+2 = (2*6 - 1) + 2 = 11 + 2 = 13. Yes. A, B and D.

OpenStudy (lilg132):

b

OpenStudy (anonymous):

lilg132 please read all of the question please :) , but b is one right answer.

OpenStudy (lilg132):

yes b a and d are correct

OpenStudy (anonymous):

as well was a

OpenStudy (lilg132):

yes b, a and d

OpenStudy (anonymous):

He said B A and D ;) Just not in alphabetical order.

OpenStudy (lilg132):

sorry A, B AND D :D